OpenStudy (anonymous):

The following table shows the probability distribution for a discrete random variable. X 24 26 27 32 35 39 P(X) 0.16 0.09 0.18 0.12 0.24 0.21 What is the mean of this discrete random variable? That is, what is E(X), the expected value of X? A. 30.5 B. 31.47 C. 32.63 D. 29.5

OpenStudy (amistre64):

\[E(x)=\sum_{i=1}^{N} x_iP(x_i)\]

OpenStudy (amistre64):

in other words, mutliply each X by its P(X), and add all the results together

OpenStudy (anonymous):

@amistre64 From where did you get this formula, and what do you mean when you say E(x)?

OpenStudy (amistre64):

E(x) is the expected value ... which is just the average, or mean, of the expectations

OpenStudy (amistre64):

spose for a game, you have 30% chance of winning 5 dollars, and have a 70% chance of losing 2 dollars the expected value of playing the game is then: 5(.3)-2(.7) = 1.5-1.4 = 10 cents on average

OpenStudy (anonymous):

10 cents what? Wining 10 cents or loosing 10 cents? Thanks for the explanation.

OpenStudy (anonymous):

but thats not one of the multiple choices

OpenStudy (amistre64):

your expected winnings over the course of many games is ... 10 cents per game that doesnt mean you will win OR lose .. but that is what your overall expectations at playing multiple games would turn out to be http://statistics.about.com/od/ProbHelpandTutorials/a/What-Is-The-Expected-Value.htm

OpenStudy (anonymous):

can you put the e(x) furmula in steps plz

OpenStudy (amistre64):

i did ... multiply each X by its P(X) and add up the results :/

OpenStudy (anonymous):

Oh okay, so you are expected to win 10 cents in each game?

OpenStudy (amistre64):

on average yes

OpenStudy (anonymous):

Oh okay thanks a lot.

OpenStudy (anonymous):

huh that link has nothing to do with my question

OpenStudy (amistre64):

yes, it does. it gives the same steps as I have mentioned to you.

OpenStudy (anonymous):

well im confused though can you do it in steps for me then

OpenStudy (amistre64):

X 24 26 27 32 35 39 P(X) 0.16 0.09 0.18 0.12 0.24 0.21 what is 24 times .16 ? what is 26 times .09 ? what is 27 times .18 ? what is 32 times .12 ? ... etc after youve multiplied all the Xs with their P(X)s ... add em up E(x) = 24(.16) + 26(.09) + 27(.18)+ ...

OpenStudy (anonymous):

ok i get it now thx

OpenStudy (amistre64):

youre welcome